OpenStudy (anonymous):

Suppose that f is a linear function with slope 2 and that f(4) = 4. Find the equation for f. (Use x as the variable.) f(x) = ?

OpenStudy (phi):

do you know the slope-intercept formula for a line ?

OpenStudy (anonymous):

change in f/change in variable ?

OpenStudy (phi):

that is the definition of slope have you seen y= mx + b ?

OpenStudy (anonymous):

oh yeah!

OpenStudy (phi):

do you remember what "m" stands for in y= mx + b ?

OpenStudy (anonymous):

slope?

OpenStudy (phi):

yes. and what is slope in your problem ? replace the "m" with the number for the slope what do you get ?

OpenStudy (anonymous):

y=2x+b

OpenStudy (phi):

next, we use f(4) = 4 I guess you have to memorize (or learn) that is a way to say when x is 4 i.e. match f(x) with f(4) to see this) then y or in this case f(x) is 4 so we put those numbers into y= 2x + b 4= 2*4 + b I would simplify 2*4 to 8 4= 8+b now add -8 to both sides to "solve for b"

OpenStudy (phi):

stuck ?

OpenStudy (anonymous):

yeah

OpenStudy (phi):

for the moment, assume (4,4) is on the line y= 2x+b can you replace x and y with their numbers? what do you get ?

OpenStudy (anonymous):

-4 = b?

OpenStudy (phi):

to get that , you should have put in 4 for y, and 4 for x and solved 4= 2*4+b 4= 8 + b 4-8 = 8-8+b -4= b so the final answer is y= 2x-4 or f(x)= 2x-4